OpenStudy (anonymous):

how do I solve : COT pi/4

OpenStudy (amistre64):

cot is the invert of tangent right? do you recall the degree version of pi/4?

OpenStudy (anonymous):

tan pi/4 = 1, so cot pi/4 is also just 1

OpenStudy (anonymous):

45 degree

OpenStudy (amistre64):

45 degrees is good :) and that has the property that sin and cos are the same value

OpenStudy (amistre64):

tan = sin/cos; which at 45 degrees is 1 right?

OpenStudy (amistre64):

and 1 flipped is just 1 again so gieskmi is good

OpenStudy (anonymous):

what is the matter with \[\frac{\pi}{4}\]? it is a perfectly good number

OpenStudy (anonymous):

Amistre64// can you solve my other questions please :3? ur explaination is good..